Cove’s Cakes is a local bakery. Price and cost information follows: Price per cake $ 13.11 Variable cost per cake Ingredients 2.35 Direct labor 1.20 Overhead (box, etc.) 0.19 Fixed cost per month $ 3,748.00 Required: 1. Calculate Cove’s new break-even point under each of the following independent scenarios: (Round your answer to the nearest whole number.)

a. Sales price increases by $1.90 per cake.

b. Fixed costs increase by $475 per month.

c. Variable costs decrease by $0.44 per cake.

d. Sales price decreases by $0.30 per cake.

2. Assume that Cove sold 430 cakes last month. Calculate the company’s degree of operating leverage. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

3. Using the degree of operating leverage calculated in Requirement 2, calculate the change in profit caused by a 6 percent increase in sales revenue. (Round your final answer to 2 decimal places (i.e. .1234 should be entered as 12.34%.))
Answer #1

1

a)

New sales 13.11+1.90 15.01
Less Variable Cost 2.35+1.20+0.19 3.74
Contribution 11.27
Total fixed Cost 3748
Contribution rate (11.27/15.01)*100 75.08%
Break even point (3748/75.08)*100 4992

b)

Sales 13.11
Less Variable Cost 3.74
Contribution margin 9.37
Total fixed cost 3748+475 4223
Contribution rate (9.37/13.11)*100 71.47%
Break even point (4223/71.47)*100 5908.77

c)

Contribution 9.37-.44 8.93
Contribution rate (8.93/13.11)*100 68.12%
Break even point (3748/68.12)*100 5502

d)

Contribution =(9.37-.30)= 9.07

Sales= (13.11-.30)=12.81

Contribution rate =(9.07/12.81)*100=70.80%

Break even =( 3748/70.80)*100=5293.78